OpenStudy (anonymous):

a 120 foot long rope is cut into 3 pieces. the first piece of rope is twice as long as the second piece of rope. the third piece of rope is three times as long as the second piece of rope. what is the length of the longest piece of rope?

OpenStudy (anonymous):

Call the peices F, S, T (first, second, third) F = 2S, T = 3S Total = F + S + T = 6S = 120 => S = ... Longest = 3S = ...

OpenStudy (anonymous):

okay

OpenStudy (gw2011):

Let x = Second Piece 2x = First Piece 3x = Third Piece The equation is: x+2x+3x=120 6x=120 x=20 ft 2x=40 ft 3x=60 ft Therefore, the third piece which 60 ft is the longest

OpenStudy (anonymous):

thank u!